OpenStudy (anonymous):

solve. 1. ln x+ln(x-2)=1 2. 3+ ln x+8 3. 3e^(-x)-4=9

OpenStudy (anonymous):

first one 1) write as a single log via \[\ln(x)+\ln(x-2)=\ln(x(x+2))\] 2) rewrite \[\ln(x(x+2))=1\] in equivalent exponential form \[x(x-2)=e\] and solve the resulting quadratic equation. you will need to use the quadratic formula

OpenStudy (anonymous):

second one is not an equation

OpenStudy (anonymous):

for number 2 it's equals 8, not plus.

OpenStudy (anonymous):

third one 1) add 4 2) divide by 3 3) write in equivalent logarithmic form 4) change the sing

OpenStudy (anonymous):

i will write it if you like, but those are the steps

OpenStudy (anonymous):

I can solve the first one for you: Rule of logs: \[loga+logb=log(ab)\] So we have: \[ln(x)(x-2)=1\] Raise both sides by e^x: \[x^2-2x=e\] You can complete the squares or use quadratic formula to get: \[1\pm\sqrt{1+e}\]

OpenStudy (anonymous):

oh in that case we do it in two steps 1) \[\ln(x)=5\] \[x=e^5\]

OpenStudy (anonymous):

For all of these the key is to realize that \[e^{ln(x)}=x\]

OpenStudy (anonymous):

@osirisis yes of course you are right completing the square is much easier because you have \[x^2-2x=e\] and -2 is even

OpenStudy (anonymous):

so @Aggie forget i mentioned the quadratic formula

OpenStudy (anonymous):

Also worth noting that ln can be used to eliminate e too. \[ln (e^x)=x\]

OpenStudy (anonymous):

osiris has written exactly what you need. also you should be able to switch from \[\log_b(x)=y\iff b^y=x\] with ease so in this case \[\ln(x)=5 \iff e^5=x\] is immediate

OpenStudy (anonymous):

So for the last porblem: \[3e^{-x}-4=9\] \[3e^{-x}=13\] \[e^{-x}=\frac{13}{3}\] \[-x=ln\frac{13}{3}\] \[x=-ln\frac{13}{3}\]

OpenStudy (anonymous):

Logs can look intimidating but if you approach them in a step by step process like normal problems they're not as bad as they seem.

OpenStudy (anonymous):

exactly! and the steps were 1) add 4 2) divide by 3 3) write in equivalent logarithmic form 4) change the sign

OpenStudy (anonymous):

so only step 3) had anything to do with exponentials and logs. the rest was elementary algebra